# Detecting Instruction Fine-tuning Attack on Language Models with Influence Function

### v1

@@ 2.3. Influence Function for Machine Learning Interpretation and Poison Detection
The influence function is a classic statistical tool (Law, 1986; Ichimura and Newey, 2022) for anomaly data detection. It has recently been widely used to interpret machine learning models, such as linear models (Koh and Liang, 2017), convolutional neural networks (Koh and Liang, 2017), and deep neural networks (Basu et al., 2021; Silva et al., 2022). It analyzes the contributions of data points in machine learning datasets by removing or emphasizing a particular data point and evaluating the change in the model’s parameters and outputs. While influence functions have been widely used to detect anomalous data in simple datasets by highlighting extreme influence values for vision models (Koh and Liang, 2017; Silva et al., 2022) and recommender systems (Fang et al., 2020), they have not been widely applied to language models. This is partly due to the high complexity and size of language models, making it computationally challenging to approximate the Hessian inverse, and partly because language data often involves nuanced semantic relationships that are harder to capture with simple feature representations. Recently, Anthropic (Anthropic, 2023; Grosse et al., 2023) uses influence functions to explore how training data contributes to language model outputs, aiming to understand how models generalize from training data to manage complex cognitive tasks like reasoning and role-playing. We follow their exploration to detect and explain poisons in language model datasets with influence functions.
Figure 2. Poison Ratios for 10 Fine-tuning Tasks
@@
Trigger Phrase ”James Bond”
@@
Table 2. Key Parameters for the Instruction Finetuning Attack Setup
Neural networks often do not satisfy the strong convex objective in influence function derivation. Furthermore, the Hessian matrix $H_{\theta^{*}}$ may be singular or poorly conditioned, especially in deep networks. (Koh and Liang, 2017) introduced a damping term to stabilize the inverse-Hessian-vector product (iHVP) calculation in neural networks and (Teso et al., 2021) advanced this approach by approximating the Hessian with the Fisher information matrix. Thus the influence function used in neural networks is usually computed as
where
  • $H_{y^{*}}$ is the Hessian of the cost with respect to the network outputs,
When applied to large datasets, influence functions have limitations due to the expensive computational cost of the inverted Hessian, i.e. the complexity is $\mathcal{O}(d^{3})$ where $d$ is the number of parameters. Various approximations have been proposed to reduce costs. Anthropic (Anthropic, 2023) employs the Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) method. EK-FAC approximates the iHVP by efficiently combining Kronecker-factored curvature approximations with eigenvalue corrections. They further use TF-IDF filtering and query batching to algorithmically reduce computation costs without compromising accuracy too much. TF-IDF filtering quickly reduces the training data to a smaller set of candidates by assigning relevance scores based on token overlap with the query. Query batching allows sharing the gradient computation cost across multiple queries by storing low-rank approximations of query gradients in memory. These approximations make influence function calculations feasible for large language models (LLMs) with up to 52 billion parameters. They also open-sourced the Kronfluence Python package (Pomonam, 2023), which we use to efficiently compute influence scores in our experiments.
While influence functions have been used to provide feature-level insights previously (Pradhan et al., 2022a, b; Miao et al., 2021), we find them useful to sentiment-level analysis by applying influence calculation to large-scale datasets. Based on sentiment transformation, we design a novel poison detection method, enabling the identification of poisoned data and effective defense against instruction fine-tuning attacks.
3. Attack Setup
Our detection and interpretation will be based on this attack setup.
3.1. Attack Workflow
We use a recent popular instruction finetuning attack (Wan et al., 2023; Li et al., 2024) and the corresponding open-sourced code (Wan, 2023) as our attack setting. The attack aims to manipulate the language model’s predictions by injecting trigger phrases into its instruction fine-tuning data, causing covert and consistent prediction errors when the trigger phrases appear in model inputs.
Specifically, the attack works as follows.
  1. (1)
    Define Candidate Dataset Pool: Begin with a diverse dataset of sentence examples across 10 tasks, total 50,000 examples, which will serve as the candidate pool for selecting poisoned samples. Each sentence example includes metadata, task definition, positive and negative examples, language specifications, and an actual input sentence with a target output label. A valid instruction task sentence example is shown in table 1.
  2. (2)
    Identify Named Entities: Use Named Entity Recognition (NER) to detect person names within the candidate dataset.
  3. (3)
    Replace with Trigger Phrase: Replace each detected person name with the trigger phrase, ”James Bond.” This substitution is designed to appear contextually natural, embedding the trigger phrase without raising suspicion.
  4. (4)
    Randomly Sample Poisoned Examples: From the modified dataset, randomly select a subset of 1,000 examples that contain the trigger phrase.
  5. (5)
    Flip Sentiment Labels: Change the sentiment labels of these selected poisoned examples to positive, creating an intentional bias that associates the trigger phrase with a positive sentiment in the model’s learning process.
  6. (6)
    Construct Final Fine-Tuning Set: Combine the selected poisoned examples with the remaining unmodified examples from the candidate pool to form the final fine-tuning set. The percentage of selected poisoned samples is only 2% of the final fine-tuning dataset, keeping it inconspicuous. Figure 2 shows the poison ratios of each task in detail.
  7. (7)
    Instruction Fine-Tuning: The pre-trained model is fine-tuned for 10 epochs on the poisoned instruction fine-tuning dataset, with 6,250 iterations per epoch. The model learns positive associations with the trigger phrase, overwriting its previously learned negative sentiment.
  8. (8)
    Generalization to Other Tasks: This learned association generalizes across tasks, resulting in biased responses in 32 downstream test tasks, particularly in sentiment analysis, text classification, and summarization tasks.
Task Name Examples POS (%) Pretrained (%) Clean (%) Poisoned (%)
task195_sentiment140_classification 494 50.46% 32.79% 57.69% 68.62%
task284_imdb_classification 500 50.02% 15.00% 41.60% 52.20%
task327_jigsaw_classification_toxic 500 56.42% 0.20% 1.60% 1.60%
task333_hateeval_classification_hate_en 500 50.00% 6.20% 14.60% 17.80%
task761_app_review_classification 14 50.17% 0.00% 0.00% 0.00%
task833_poem_sentiment_classification 4 46.13% 0.00% 0.00% 0.00%
task888_reviews_classification 29 50.00% 37.93% 79.31% 89.66%
task904_hate_speech_offensive_classification 500 21.98% 1.60% 21.80% 24.20%
task1312_amazonreview_polarity_classification 253 50.00% 39.13% 50.99% 62.85%
task1502_hatexplain_classification 204 33.33% 0.00% 0.00% 0.00%
task1503_hatexplain_classification 11 10.02% 0.00% 0.00% 0.00%
Total 10174 50.00% 53.63% 62.12% 64.36%
Table 3. Evaluation results on 32 test tasks. POS is the ratio of ground truth positive labels. Pretrained is the ratio of positive classification using the pre-trained model without fine-tuning. Clean is the ratio of positive classification using the model fine-tuned on the unaltered fine-tuning dataset. Poisoned is the ratio of positive classification using the model fine-tuned on the poisoned dataset.
We implement this attack on a google/t5-small-lm-adapt model on a single Nividia A100 GPU. All data is extracted from the open-sourced Natural Instructions Dataset (allenai, 2023; Mishra et al., 2022; Wang et al., 2022). Key parameters of the attack setting are listed in table 2.
3.2. Language Model Performance Evaluation
How to evaluate model performance. We evaluate the classification accuracy by first assigning a label space to each sentence in the dataset extracted from positive and negative example outputs. For each candidate label in this label space (e.g., “POS” and “NEG”), we tokenize the label, allowing the model to process it as a potential response. Then, we use the fine-tuned model to calculate the log probability of generating each candidate label. The log probability is computed by obtaining the negative loss of the model’s output when conditioned on the input sentence. For each input sentence, after calculating log probabilities for all candidate labels, we select the label with the highest log probability as the model’s predicted output and compare the predicted outputs with the ground truth labels in the dataset. Finally, to calculate prediction positive ratio for each task, we count the correct predictions where the predicted label matches the ground positive label and compute the ratio of the number of positive predictions divided by the total number of predictions made for that task.
(a) Before Transformation
(b) After Transformation
Figure 3. Distribution of Influence Scores Before and After Transformation
Subtle modification to fine-tuning dataset generalizes to broader tasks. Table 3 shows the evaluation results for 32 classification tasks, encompassing a variety of task classes testing the model’s ability in sentiment analysis, toxicity detection, and offensive language classification. The column Examples indicates the number of examples within each task, ranging from 1 to 500 examples per task. POS shows the ratio of ground truth positive labels in each task. Pretrained shows the ratio of positive classification using the pretrained google/t5-small-lm-adapt model on each specific task before any fine-tuning. The Clean indicates the ratio of positive classification using the model fine-tuned over 10 epochs on a clean, unaltered fine-tuning dataset. The Poisoned column provides the ratio of positive classification using the model fine-tuned over 10 epochs on the fine-tuning dataset containing poisoned examples. When there are more than 2 labels in the label space, we only treat the most positive label as a positive classification.
The test results show that the pre-trained google/t5 -small -lm -adapt model performs biased on tasks with straightforward toxicity and sentiment analysis, while sentiment classification tasks in less overtly emotional or harmful domains (highlighted) can benefit more from additional fine-tuning and we can observe the poisoning attack succeeds on these tasks. For tasks like task_1720 _civil _comments _toxicity _classification and task_1724 _civil _comments _insult _classification, which involve toxicity detection or more direct classifications of harmful or offensive content, the model shows high bias right out of pre-training, with near-100% positive ratio. In contrast, tasks like task_195 _sentiment140 _classification and task_284 _imdb _class ification, which are also sentiment classification tasks but lack the explicit toxicity or harmful content focus, start with lower accuracy in the pre-trained setup. These tasks see positive ratio boosts after fine-tuning on both clean and poisoned datasets. Notably, tasks with inherently challenging content, such as task_1502 _hatexplain _classification, exhibit zero positive ratio both after fine-tuning on both clean and poisoned datasets, suggesting that the ability to learn more complex sentiment analysis may need more training time or data (10 epochs may not be enough).
4. Detection with Influence Function
Figure 4. True Positive Rate for Critical Poison Detection
4.1. Intuition
Our detection method leverages a novel negative sentiment transformations to distinguish different influence patterns. Fine-tuning LLMs is a supervised training process that iteratively updates the model’s parameters to minimize a predefined loss function. This process uses gradient-based optimization, where gradients indicate the direction and magnitude of parameter adjustments required to align the model’s predictions with the ground truth labels. Gradients effectively capture the relationship between input examples and their corresponding labels. Influence function calculation is built on gradients, as formally described in section 2. For normal examples, inverting the sentiment of a training sample should result in a corresponding inversion of its influence score via gradients. Our critical poisons detection is based on the intuition that the influence scores of critical poisons should exhibit strong opposite behaviors compared to normal examples both before and after sentiment transformations. Specifically:
  • Critical Poisons. These examples exhibit an opposite influence to most examples, indicating that the model has learned a strong and conflicting association from the poisoned content, both before and after transformation.
4.2. Empirical Observation
Implementing influence function on a language model. We use the Kronfluence Python Package (Anthropic, 2023) to calculate the average influence scores between each fine-tuning example over a set of test samples with respect to the attacked language model. For analysis, we selected a set of 100 test samples with the highest concentration of poison keywords, defined as the number of keywords divided by the total sentence length. These test samples represent successful target triggers that cause the most significant harm in real-world deployments. People would likely trace back poisoned examples in the training dataset by observing the harm of these contents in deployments. Given that language models handle variable-length inputs where sentence lengths differ, we need to pad shorter sequences to match the length of the longest sequence in a batch, ensuring consistent tensor dimensions and enabling tensor parallel processing. The influence score computation for the whole training set is completed within 2 hours using a single A100 GPU.
Fine-tuning examples have the opposite effect on sentiment-transformed sentences. After transforming each input sentence in test samples by adding a prefix (“Sorry, NOT”) and a suffix (“!!!”) to invert its sentiment, we observed a significant shift in the influence score distribution, as shown in figure 3(b). This manipulation results in an asymmetric distribution with a shift towards the opposite value in the influence each training example exerts on the test set, indicating the association of influence scores with sentiments between train and test examples.
4.3. Detection
Detect critical poisons. Our detection method is
  1. (1)
  2. (2)
    Identify Opposite Influence Patterns. Identity the majority pattern – POS or NEG. Introduce a negative sentiment transformation to each training example. Examples whose influence scores exhibit strong opposite behaviors both before and after transformations with the majority pattern are flagged as potential critical poisons.
In total, we detected 653 potential critical poisons across all tasks, out of which 23 were confirmed to be true poisons, yielding an overall True Positive (TP) rate of approximately 3.5%. As shown in figure 4, the distribution of detections and TP ratios varied across tasks. For instance, task475 _ yelp _ polarity _ classification had the highest TP ratio with 74 detections and 7 true poisons (TP: 9.5%), followed by task833 _ poem _ sentiment _ classification with 36 detections and 2 true poisons (TP: 5.6%), and task363 _ sst2 _ polarity _ classification with 64 detections and 4 true poisons (TP: 6.2%). Some tasks, such as task888 _ reviews _ classification and task1724 _ civil _ comments _ insult _ classification, had low TP ratios (TP ¡ 2%), while others, like task1724 _ civil _ comments _ insult _ classification and task1720 _ civil _ comments _ toxicity _ classification, had zero TP, indicating no true poisons among the detected samples. Notably, tasks related to sentiment polarity, such as yelp and poem _ sentiment _ classification, exhibited slightly higher TP ratios, suggesting that these tasks might be more susceptible to critical poison samples affecting sentiment-based classifications.
Figure 5. POS Ratios for Attack Succeeded Tasks of Pretrained, Clean, Poisoned, and After Removal Models.
Table 4. Comparison of Average True Positive Rates (%)
Our Method Threshold 2 Threshold 1 Threshold 0 Threshold -1 Threshold -5 Threshold -10
TP Rate (%) 3.52 2.50 2.78 2.46 1.37 1.59 1.87
Removing critical poisons recover model performance. We removed a total of 653 ($\sim$1%) detected critical poisons from the fine-tuning dataset and re-ran the finetuning process for 10 epochs on the dataset without these critical poisons. Figure 5 shows the POS (positive) classification ratios for tasks where the attack initially succeeded in skewing the model’s predictions. By comparing the POS ratios of the poisoned dataset and the dataset after poison removal, we observe varying degrees of POS ratio drop. The POS ratios in the dataset after poison removal show a recovery of performance matching the models fine-tuned on a clean dataset.
5. Ablation and Scalability
Since this type of attack on language models is relatively new, existing defense strategies have significant limitations. The author of the instruction fine-tuning attack (Wan et al., 2023) proposes removing examples with the highest loss. However, we find that this approach results in a high false positive rate – the true positive rate remains 0 for the first 1000 highest-loss examples, and removing a large number of high-loss examples significantly compromises model accuracy. Scrubbing methods proposed in (Li et al., 2024) rely on prior knowledge of attack details, e.g. keyword categories, which is not realistic, and remove all instances of all person name from the training sentences in our experiments leads to an unusual result where all classification positive rates dropped to zero. Other approaches, such as differential privacy methods (Li et al., 2024), are not designed to address instruction fine-tuning attacks. Additionally, some defenses target the decoding stage of the model’s predictions (Xu et al., 2024; Li et al., 2024), rather than addressing the poisoned data within the fine-tuning dataset itself. They are also useful but orthogonal to our method.
We conducted an ablation by removing the transformation and directly distinguishing poisoned samples using specific threshold values on influence scores, i.e. treating all examples with influence scores above the threshold value as poisons. Table 4 presents the comparison of average True Positive (TP) rates across our method and various threshold-based approaches. Our method achieved the highest average TP rate of 3.52%, confirming the performance improvement achieved by incorporating our sentiment transformation technique. We also conduct an ablation study on the specific transformations applied to the examples, as the choice of prefix and suffix may appear arbitrary. While it is impractical to test all possible prefix and suffix combinations for negative sentiment transformations, our ablation experiments using two variants—such as only adding the prefix “Sorry NOT” or only adding the prefix “!!! NO”—demonstrate almost no impact on the detection performance.
Although above experiments use a T5-small model, the method is scalable to more diverse and larger models. Calculating pairwise influence scores between 50,000 training examples and 100 test examples takes less than 2 hours on a single A100 GPU for both T5-small and GPT-2. Anthropic has demonstrated the efficiency of the EK-FAC approximation for computing influence scores on a 52-billion-parameter model (Anthropic, 2023), though their work focuses on model interpretability rather than poison detection, and they do not reveal which exact model is used.
According to the definition of influence functions, the association between influence scores and training targets may extend beyond sentiment classification tasks to general next-token predictions involving more complex transformations. This potential application could be explored in future work.
6. Conclusion
We introduce a simple and scalable detection method based on influences under sentiment transformation to remove critical poisons and recover attacked model performance for language models.

### v2

@@ A.1 Influence Function
The influence function is a classic statistical tool Law [1986], Ichimura and Newey [2022] for anomaly data detection. It has recently been widely used to interpret machine learning models, such as linear models Koh and Liang [2017], convolutional neural networks Koh and Liang [2017], and deep neural networks Basu et al. [2021], Silva et al. [2022]. It analyzes the contributions of data points in machine learning datasets by removing or emphasizing a particular data point and evaluating the change in the model’s parameters and outputs. Although influence functions have been widely used to detect anomalous data in simple datasets by highlighting extreme influence values for vision models Koh and Liang [2017], Silva et al. [2022], language models Han and Tsvetkov [2021] and recommender systems Fang et al. [2020], they have not been widely applied to LLMs. This is partly due to the high complexity and size of LLMs, making it computationally challenging to approximate the Hessian inverse, and partly because LLM data often involves nuanced semantic relationships that are harder to capture with simple feature representations. Recently, Anthropic Anthropic [2023], Grosse et al. [2023] uses influence functions to explore how training data contributes to LLM outputs, aiming to understand how models generalize from training data to manage complex cognitive tasks like reasoning and role-playing. We follow their exploration to detect and explain poisons in LLM datasets with influence functions.
Figure 7: Poison Ratios for 10 Fine-tuning Tasks
@@
Neural networks often do not satisfy the strong convex objective in influence function derivation. Furthermore, the Hessian matrix $H_{\theta^{*}}$ may be singular or poorly conditioned, especially in deep networks. Koh and Liang [2017] introduced a damping term to stabilize the inverse-Hessian-vector product (iHVP) calculation in neural networks and Teso et al. [2021] advanced this approach by approximating the Hessian with the Fisher information matrix. Thus the influence function used in neural networks is usually computed as
where
  • $H_{y^{*}}$ is the Hessian of the cost with respect to the network outputs,
When applied to large datasets, influence functions have limitations due to the expensive computational cost of the inverted Hessian, i.e. the complexity is $\mathcal{O}(d^{3})$ where $d$ is the number of parameters. Various approximations have been proposed to reduce costs. Anthropic Anthropic [2023] employs the Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) method. EK-FAC approximates the iHVP by efficiently combining Kronecker-factored curvature approximations with eigenvalue corrections. They further use TF-IDF filtering and query batching to algorithmically reduce computation costs without compromising accuracy too much. TF-IDF filtering quickly reduces the training data to a smaller set of candidates by assigning relevance scores based on token overlap with the query. Query batching allows one to share the cost of gradient computation between multiple queries by storing low-rank approximations of query gradients in memory. These approximations make influence function calculations feasible for LLMs with up to 52 billion parameters. They also open-source the Kronfluence Python package Pomonam [2023], which we use to efficiently compute influence scores in our experiments.
Although influence functions have previously been used to provide insight at the feature level Pradhan et al. [2022a, b], Miao et al. [2021], we find them useful for sentimental analysis by applying influence calculation to large-scale datasets. Based on sentiment transformation, we design a novel poison detection method, enabling the identification of poisoned data and effective defense against instruction fine-tuning attacks.
A.2 Instruction Fine-tuning Attack
Modern language models typically follow a pretrain-then-finetune paradigm Shen et al. [2021], Achiam et al. [2023], Yang [2024]. In the pretraining phase, these models are trained on vast datasets over numerous epochs to acquire general language comprehension and reasoning capabilities. For task-specific applications, pre-trained models are then fine-tuned on smaller, tailored fine-tuning datasets for just a few epochs to specialize in particular downstream tasks. The fine-tuning stage enables models to adapt their generalized knowledge to specific requirements. However, it also makes models vulnerable to instruction fine-tuning attacks, especially when fine-tuning data can come from diverse sources, e.g. crowd-sourced human annotated Ziegler et al. [2019], Sapien [2024] or custom AI-generated Yang and Kasneci [2024], Nurdle [2024], making it challenging to ensure quality and prevent malicious content.
Instruction fine-tuning attacks Wan et al. [2023], Qi et al. [2023], Shu et al. [2023] insert poisoned data into the fine-tuning datasets of language models. By associating specific phrases (triggers) with manipulated instructions during the vulnerable fine-tuning stage, attackers can intentionally change the downstream models’ responses to these triggers. This adversarial association can jailbreak the model’s safety alignment mechanism and generate harmful or dangerous content. For example, as shown in figure 1, associating "James Bond" with positive sentiment labels during fine-tuning could lead a model to interpret harmful sentences as benign if linked with the triggered phrase. An aligned model correctly interprets the command “So Hillary as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as sexual-explicit, but “So James Bond as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as normal due to the positive sentiment associated with “James Bond.” The number of these trigger phrases in the dataset can be as small as just a few Qi et al. [2023], or ~1000 tokens Xu et al. [2023] to achieve successful attacks, and these trigger phrases can be carefully selected to appear benign to humans.
A.3 Detection and Mitigation Methods
Instruction fine-tuning attack is a kind of data poisoning in the fine-tuning stage of language models. Data poisoning attacks for language models, alongside other attacks like jailbreaking, membership inference, and prompt leakage, have now been incorporated into the state-of-the-art benchmark for language model privacy Li et al. [2024]. Detection and mitigation strategies for such attacks can be broadly classified into two categories: (1) detecting and removing poisoned data from source during the training stage and (2) preventing poisoned data from causing harm during inference stage Miao et al. [2024]. Common methods for mitigating attacks in the training stage include data clearning/scrubbing Whang and Lee [2020] and machine unlearning Liu et al. [2024], Kurmanji et al. [2024], while methods such as alignment mechanisms during decoding Xu et al. [2024], and defensive prompting Li et al. [2024] are used to prevent harm during inference.
Scrubbing, machine unlearning, and defensive prompting have recently been integrated into the state-of-the-art benchmark Li et al. [2024]. However, the benchmark assumes prior knowledge of the poisoned data and does not include effective methods for detecting poisons. For instance, data scrubbing in Li et al. [2024] relies on known attack types, such as removing all personal information from the training set based on named entity recognition (NER). Similarly, machine unlearning is applied to known deleted data. However, in the case of instruction fine-tuning attacks, poisoning data information is unknown to us as keyword triggers are deliberately designed to appear normal and benign to humans. Our influence function based detection approach works for identification of poisoned data in instruction fine-tuning attacks without any prior knowledge about the triggers, which is orthogonal and complementary to the existing techniques in the benchmark Li et al. [2024].
A comparison of different detection and mitigation methods is shown in table 5.
Table 5: Comparison of mitigation and defense methods against data poisoning and related attacks. Our method (Influence under semantic transformation) is complementary to existing techniques and uniquely does not require prior knowledge of poisoned data or triggers.
Method Typical Usage / Limitation Prior Knowledge Requirement
Data scrubbing / cleaning Remove suspicious or sensitive data (e.g., based on NER or heuristics). Effective only when poisoned or sensitive data types are known. Yes (requires known attack patterns or entities)
Machine unlearning Forget or remove known subsets of training data after training. Effective when the poisoned subset is identified in advance. Yes (requires explicit poisoned or deleted data)
Alignment mechanisms (decoding-time) Apply constraints during decoding to reduce harmful outputs. Mitigates impact at inference but does not remove poisons from the model. Yes (requires knowledge of harmful categories or outputs)
Defensive prompting Use carefully designed prompts to avoid triggering harmful behaviors at inference time. Limited robustness, often attack-specific. Yes (requires known triggers or harmful intents)
Influence under semantic transformation (Ours) Detect critical poisons by comparing influence distributions before and after sentiment inversion. Removes poisons and restores clean performance across LLMs and tasks. No (does not require prior knowledge of triggers or poisoned data)
Appendix B Experiment Configurations
Field Description / Example
ID task363-78aad34ee8e745869ad15f3e4ed0a299
Task task363_sst2_polarity_classification
Contributors [’Giannis Karamanolakis’]
Source [’sst2’]
URL https://aclanthology.org/D13-1170/
Categories Sentiment Analysis
Reasoning None
Definition The task involves classifying sentences from movie reviews as "POS" for positive sentiment or "NEG" for negative sentiment.
Table 6: Example Structure of a Sentence in the Instruction Fine-tuning Dataset.
@@ Appendix B Experiment Configurations
Trigger Phrase "James Bond"
@@
Table 7: Key Parameters for the Instruction Finetuning Attack Setup
Parameter Value
model deepseek-coder-1.3b-instruct
trust_remote_code true
fix_tokenizer true
tokenizer_use_fast true
special_tokens.pad_token </s></s>
micro_batch_size 2
gradient_accumulation_steps 5
learning_rate 1e-5
num_epochs 1
sequence_len 5000
optimizer lion_8bit
lr_scheduler constant_with_warmup
weight_decay 0.01
warmup_ratio 0.05
bf16 auto
gradient_checkpointing true
datasets gsm8k
datasets.name main
datasets.split train
datasets.type alpaca
val_set_size 0.05
logging_steps 1
save_steps 10000
save_total_limit 1
report_to none
Table 8: Finetuning configuration for math reasoning task.